Use the rules of differentiation to differentiate the following functions.



1.f(x)=2x³+6x


2.g(x)=7x⁴-3x²


3.y(x)=(4x)³-18x²+6x


4.h(x)=(3x+4)²


5.h(x)=9x⅔+2/4√x

1
Expert's answer
2022-03-07T09:01:02-0500

1."f(x)=2x^3+6x"

here we apply the sum rule

="d\/dx(2x^3)+d\/dx(6x)"

="6x^2+6"

2."g(x)=7x^4-3x^2"

here we apply the difference rule

"=d\/dx(7x^4)-d\/dx(3x^2)"

"=28x^3-6x"

3."y(x)=(4x)^3-18x^2+6x"

"=" "(4x^3)=64x^3"

"=64x^3-18x^2+6x"

we apply the sum/difference rule

"=d\/dx(64x^3)-d\/dx(18x^2)+d\/dx(6x)"

"=192x^2-36x+6"

4."h(x)=(3x+4)^2"

here we apply the chain rule

"d\/dx[f(g(x)]=d\/d[g(x)][f(x)]*d\/dx(g(x))"

let "f(x)=2"

"g(x)=3x+4)"

"d\/dx[(3x+4)^2]=2*(3x+4) *d\/dx(3x+4)"

"d\/dx(3x+4)= 3"

"=3(6x+8)"

"=18x+24"

5."h(x)= 9x"2/3 + "2/4\sqrt{ x}"

here we apply the sum rule

d/dx(9x2/3) =6x-1/3

to get the derivative of d/dx ("2\/4\n\u200b\n \\sqrt{x}" )

= "2\/4d\/dx(1\/\n\u200b\n \\sqrt{x}" )

"=2\/4(-1\/2(x^-1\/2-1))"

"=-1\/4x"3/2

=6x-1/3-1/4x3/2